OpenStudy (anonymous):

please help me confirm my answer because I am bad at math! so I have the problem (-1/2x^5)(-1/9x) and I got the answer of 2/9x^6, is this correct?

OpenStudy (anonymous):

should be 1/18x^6

OpenStudy (anonymous):

no its suppose to be 1/8 x^6

OpenStudy (anonymous):

man I suck at this, do either or both of you mind showing me your work so that I can know how you got the answer?

OpenStudy (anonymous):

reason is because you simply multiply common factors together in this case just multiply -1/2 * -1/9 = 1/18 other common: x^5 * x = x^6

OpenStudy (anonymous):

multiply the numerators (-1) * (-1) = 1 then the denominators (2x^5)(9x) = 18x^6

OpenStudy (anonymous):

oh I thought I had to have the same denomitor to multiply, so that's where I went wrong, thanks!

OpenStudy (anonymous):

denominator I mean

OpenStudy (anonymous):

Nope that's for addition.